Assume your PIN is a 2-digit number, in which each digit is a number from 1 to 9, and numbers can be used only once. What is the probability that a thief who steals your debit card will guess your PIN on the 10th attempt after failing on the previous nine attempts? Assume the thief is reasonably intelligent and only tries a PIN number once.
Answer #1

Total number of 2 digit distinct pin combinations possible = 9*8 = 72

As the thief enters wrong pin for first 9 attempts , he will be left with only 63 distinct pin combinations including the correct pin.

So the probability of guessing correct pin on 10th attempt

= 1/63